X has a molar mass of 46.31, what is the percentage composition by mass of X in X3PO4?

 


1
Expert's answer
2021-06-14T13:37:07-0400

A molar mass of X3PO4 equals:

Mr(X3PO4) = (3 × 46.31 g/mol) + 31 g/mol + (3 × 16 g/mol) = 217.93 g/mol

The percentage composition by mass of X in X3PO4 equals:

w % = [Mr(X) / Mr(X3PO4)] × 100 % = [46.31 g/mol / 217.93 g/mol] × 100 % = 21.25 %


Answer: 21.25 %